OpenStudy (anonymous):

How would I graph x-y=5

OpenStudy (amistre64):

find to points the fit; and connect the dots

OpenStudy (amistre64):

two points that is ....

OpenStudy (anonymous):

I'm trying to see if I graphed it write on paper.

OpenStudy (amistre64):

how did you graph it?

OpenStudy (amistre64):

when x=0, y=? (0,y) x-y = 5 0-y = 5 y = -5 (0,-5) when y=0, x=? (x,0) x-y = 5 x-0 = 5 x = 5 (5,0)

OpenStudy (amistre64):



OpenStudy (anonymous):

okay. i was thinking I had to graph more points. thanks !!

OpenStudy (anonymous):

Hey, I got the same graph for y+5=0. is this correct?

OpenStudy (amistre64):

y+5 = 0 -5 -5 -------- y = -5 y = -5 is just a flat line thru y=-5